Circular No. 6727

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 F. E. Marshall, Goddard Space Flight Center (GSFC); and J. K. Can-

nizzo and R. H. D. Corbet, GSFC and Universities Space Research Association, on behalf of the RXTE Science Operations Team, report the detection with the RXTE PCA of an x-ray source whose position appears consistent with that reported on IAUC 6726: “Multiple slews across this position, beginning at Aug. 28.890 UT and lasting a total of 40 min, revealed a source intensity in the band 2–10 keV of ∼ 0.5 mCrab. Since we know of no catalogued x-ray sources at this position, we believe this is the detection of an x-ray afterglow from this burst. A refined error box should be possible after more detailed analysis.”

SUNGRAZING COMETS C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discoveries by Doug Biesecker and Bernard McCarty (University of Birmingham) of four more apparent Kreutz-sungrazing comets in archival SOHO-LASCO C3 coronagraph data. Measurements by St. Cyr and Biesecker have been reduced by G. V. Williams and are given on the MPECs cited below. No tails were seen in the images of C/1996 O2 (MPEC 1997-Q04) and C/1996 O3 (MPEC 1997-Q05), whose peak apparent magnitudes were ∼ 8. Tails were evident for comets C/1996 O1 (MPEC 1997-Q03) and C/1996 O4 (MPEC 1997-Q06), whose peak magnitudes were estimated as ∼ 7 and 6, respectively. Approximate discovery positions are as follows:

1996 UT  α_{2000}  δ_{2000}  Comet
July 21.419  7^{h}50.^{m}1  +18^{o}18′  C/1996 O1
23.195  7 51.4  +17 16  C/1996 O2
25.120  8 02.1  +17 33  C/1996 O3
27.031  8 06.0  +16 28  C/1996 O4

COMET 43P/WOLF-HARRINGTON Total CCD magnitude estimates: Aug. 8.07 UT, 14.1 (A. Galád and

A. Pravda, Modra, Slovakia, 0.6-m reflector); 12.10, 14.1 (H. Mikuž, Crni Vrh Observatory, Slovenia, 0.36-m reflector + V filter); 22.08, 13.3 (Galád and Pravda); 26.12, 14.3 (Galád and Pravda).

1997 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green